OpenStudy (ghdoru11):

Find the area of the plane domain bounded by the curves: y = 3 √x2, y = 1 4+x2, x = 2 and x = 2√3

OpenStudy (ghdoru11):

\[y=\sqrt[3]{x^2} , y=\frac{ 1 }{ 4+x^2 } , x=2 , x=2\sqrt{3}\]

OpenStudy (loser66):

\[\large \int_2{^{2\sqrt{3}}}~~ x^{2/3}-\dfrac{1}{4+x^2}dx\]

OpenStudy (ghdoru11):

ty

OpenStudy (ghdoru11):

can you write the step by step solve?

OpenStudy (loser66):

those are steps!! 1) draw the picture out 2) set up the integral 3) take integral . I let you do this step 4) done. the first step is the most important one to know how to set up the integral. If you skip this step, you have to set up the integral algebraic way and it is not a piece of cake. However, it has an outlet to go, though.

OpenStudy (ghdoru11):

thank you